OpenStudy (anonymous):

Solve each system of equations. Your answer will be in the (x,y) order. One equation and plug it into the other equation, then solve for x. Once you know x, then solve for y. y = x -1 y = 3x – 7 So I'm going to take the first equation and substitute/plug it into the second equation. x - 1 = 3x – 7 Now I need to get the letters on one side and the numbers on the other side. subtract x over to the right side -1 = 2x – 7 add 7 over to the left side 6 = 2x solve for x by dividing by 2 3 = x Now that we know x, we can plug 3 in for x into either equation and solve for y. So y = x - 1, y = 3-1, so

OpenStudy (agent0smith):

That looks correct.

OpenStudy (anonymous):

okay here is the Q 1.) y = 2x y = 12 – x A.) (4,8) B.)(1,2) C.) (3,2) 2.) y = 2x - 4 y = x – 1 A.) (4,8) B.)(1,2) C.) (3,2) 3.) y = 3x - 1 y = x + 1 A.) (4,8) B.)(1,2) C.) (3,2)

OpenStudy (agent0smith):

Try your method again on questions 1 and 2 :)

OpenStudy (anonymous):

i still need help my mom helped me with at....

OpenStudy (agent0smith):

for this one: y = 2x y = 12 – x First put them equal to each other: 2x = 12-x then try to find x. Then use it to find y.

OpenStudy (anonymous):

(3,2)?

OpenStudy (anonymous):

yup! its (3,2)

OpenStudy (anonymous):

2.) y = 2x - 4 y = x – 1 A.) (4,8) B.)(1,2) C.) (3,2)

OpenStudy (anonymous):

(3,2)

OpenStudy (anonymous):

3.) y = 3x - 1 y = x + 1 A.) (4,8) B.)(1,2) C.) (3,2) ((:

OpenStudy (agent0smith):

2x = 12-x this one isn't gonna give x=3